# TransMamba: Flexibly Switching between Transformer and Mamba

### v1

@@ Abstract
Transformers are the cornerstone of modern large language models, but their quadratic computational complexity limits efficiency in long-sequence processing. Recent advancements in Mamba, a state space model (SSM) with linear complexity, offer promising efficiency gains but suffer from unstable contextual learning and multitask generalization. This paper proposes TransMamba, a novel framework that unifies Transformer and Mamba through shared parameter matrices (e.g., QKV and CBx), and thus could dynamically switch between attention and SSM mechanisms at different token lengths and layers. We design the Memory converter to bridge Transformer and Mamba by converting attention outputs into SSM-compatible states, ensuring seamless information flow at TransPoints where the transformation happens. The TransPoint scheduling is also thoroughly explored for further improvements. We conducted extensive experiments demonstrating that TransMamba achieves superior training efficiency and performance compared to baselines, and validated the deeper consistency between Transformer and Mamba paradigms, offering a scalable solution for next-generation sequence modeling.
@@ 1 Introduction
Transformers (Vaswani et al., 2017; Achiam et al., 2023; Touvron et al., 2023) are the foundation and mainstream model of modern deep learning (Zhao et al., 2023), showing dominating power in language modeling. Recently, Mamba has emerged (Gu & Dao, 2023) and been verified in various fields. Compared with Transformer, Mamba has linear computational complexity, high efficiency in processing long sequences, and lower training and inference costs (Qu et al., 2024). Nevertheless, its contextual learning and multi-task generalization capabilities are unstable (Waleffe et al., 2024). Transformer and Mamba have their own strengths and complement each other.
However, Transformer and Mamba have their own flaws that cannot be addressed by naive layer-shared hybrid structures (Yuan et al., 2024; Yang et al., 2024). For example, Transformer has faster training for short contexts while Mamba has better efficiency in longer contexts (see Table 2). Moreover, the naive static hybrid model has structural restrictions such as the order of Mamba and Transformer, mandatory ratios, etc (Lieber et al., 2024; Dao & Gu, 2024). The performance of the Hybrid model will deteriorate if these specific rules are not met, which greatly limits the exploration and breakthrough of the model.
Recently, Mamba2 (Dao & Gu, 2024) further enhances the performance of Mamba series, which reveals the surprising consistency of the attention of Transformer and the State Space Model (SSM) of Mamba. Furthermore, (Wang et al., 2024) performed distillation between Mamba and Transformer, distilling the QKV parameters of Attention to obtain CBx of SSM, verifying that the parameters can be interactively transferred as shown in Table 2. These motivate us that we can bravely utilize a set of shared parameters of QKV and CBx to build a joint Transformer-Mamba framework, which could flexibly decide which structure is suitable for the current training/inference in different layers/token lengths, taking advantage of both structures to balance effectiveness and efficiency while ensuring structural flexibility. Intuitively, to obtain the efficiency advantages of both structures, we can make the model adopt the Transformer mechanism for training on relatively short contexts and the SSM mechanism on long contexts. As shown in Figure 3, such prototype framework has only one set of parameters to flexibly switch between Transformer and Mamba for LM. In the first N tokens of the sequence, the parameter matrix is calculated using the attention mechanism. At a specific node in the sequence (which we call TransPoints from Transformer mode to Mamba2 mode), the parameter matrix is converted to the SSM mechanism for subsequent sequence generation, so as to achieve better training efficiency with better performance in sequences of different lengths.
@@ 2.1.1 Basic Notions and Consistency of Attention and SSM
We use the classic notation from the Transformer and Mamba papers. $\mathbf{QKV}$ denotes the key parameters (query, key, value) of Attention, and $\mathbf{L}$ denotes the additional mask matrix. $\mathbf{CBx}$ represents the key parameters in the SSM, $\Delta$ is used to control the discrete step size in SSM, and $\mathbf{A}$ is used to describe the global dependencies of the hidden state, which is similar to the mask matrix in Attention (Gu & Dao, 2023). In order to satisfy the classic symbolic representation and clear expression, we use $\mathbf{H}$ to denote the input embeddings for attention and SSM. The corresponding calculations are shown in Table 2.
Mamba2 (Dao & Gu, 2024) compares the underlying mechanisms of Transformer and Mamba, and introduces the dual form of SSM to illustrate the consistency between the two. In Table 2, we can find that the core mechanisms of Transformer and Mamba (attention and SSM) are completely symmetrical. Wang et al. (2024) aligned the QKV of the transformer weights with CBx of Mamba and performed distillation, achieving improved results on chat and long-text benchmarks. This once again shows that the core weights of Transformer and Mamba are transferable and unified. The above theories and research inspired us to build a bolder framework of TransMamba with a unified architecture of Transformer and Mamba.
@@ 2.1.2 Efficiency of Attention and SSM with different token lengths
As shown in Table 2, recent work (Dao & Gu, 2024) theoretically summarizes the FLOPs of Attention and SSM. $\mathrm{T}$ denotes the sequence length, $\mathrm{N}$ denotes the state dimension and $\mathrm{P}$ denotes the TransPoint value. When T is greater than N, Transformer has an advantage in efficiency on shorter sequences, while Mamba is efficient at training on long sequences due to its linear complexity of T. This advantage of Mamba’s training efficiency on longer contexts is present with most of the commonly-used model sizes, which also forms the motivation of our TransMamba that attempts to unleash the maximum potential of the flexible hybrid Transformer-Mamba structure in terms of effectiveness and efficiency.
@@ 2.4.2 Detailed TransPoint Schedule Designing
TransPoint schedule from token length aspect. Due to the flexibility of TransPoint scheduling at different layers and token lengths, the model structure of TransMamba varies at different positions. Suppose the number of layers of the model is ${L}$ and the length of the sequence is ${T}$, there are ${L}*{T}$ possible TransPoint schedules for our TransMamba with fixed parameters. We denote the value of TransPoint as ${P}$ (indicating that the tokens before position ${P}$ are modeled via Transformer and those after ${P}$ are encoded via Mamba), and we have the FLOPs for a TransMamba layer as follows:
@@
TransPoint schedule from layer aspect. We find that simply setting a global Transpoint for all layers (e.g., simultaneously at length 4,096) will result in unsatisfactory performance due to the sudden switching. To achieve better results, we need to set more diverse TransPoints at the layer level, which could gradually guide the model structure from pure Transformer to Mamba differently at various layers. To enable a smoother transformation, the TransPoints are placed separately but as close as possible to the optimal efficient $P$ according to Eq. 10. Specifically, our TransPoints cycle is performed every 8 layers, referring to the work (Dao & Gu, 2024) on hybrid structure. The mean of TransPoints is set slightly smaller than the optimal efficient point to enable more Mamba layers for better performance. TransPoints gradually transition from the beginning to the end of the sequence in a logarithmic trend (i.e., 0, 128, 256, 512, 1024, 2048, 4096, 8192), ensuring dispersion and smoothness.
@@ 3.1 Experiment Setup
We developed three baseline model families with various sizes (400M, 1.5B): Transformer (Shoeybi et al., 2019), Mamba2 (Dao & Gu, 2024), and Hybrid (Lieber et al., 2024). All models are developed based on the Megatron-LM library. Specifically, all models are unified in model size for fair comparisons. The models are pre-trained utilizing collected in-house dataset which consists of a cleaned combination of Chinese and English datasets. We trained all models on 83 billion tokens for all models. For evaluation, we aim to achieve robust conclusions across diverse domains. We conducted comprehensive evaluations involving 8 English tasks, including ARC-E, ARC-C (Clark et al., 2018), CoQA (Reddy et al., 2019), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), PhoneBook (Waleffe et al., 2024), BoolQ (Clark et al., 2019), LongBench-v2 (Bai et al., 2024). More details are shown in Appendix A.3.1 and A.3.2.
@@ 4 Related Works
Transformer has always been the focus of language model research (Beltagy et al., 2020; Liu et al., 2021; Tang et al., 2024), but its limitations in processing long sequences (Zhou et al., 2021; Behrouz et al., 2024) and the memory pressure caused by KV cache (Wang et al., 2020; Dao et al., 2022) are also difficult to solve. Mamba has the advantage of linear complexity based on the state space model (Gu & Dao, 2023; Zhang et al., 2024), but struggles in modeling complex contexts (Xiao et al., 2024).
Hybrid models (Chen et al., 2024; Lou et al., 2024; Ren et al., 2025) that combine the two are emerging, but most of the work simply cascades them (Hatamizadeh & Kautz, 2024; Lieber et al., 2024). Recent work (Dao & Gu, 2024; Han et al., 2024; Wang et al., 2024) has revealed the consistency of the underlying mathematics between them. However, there is no work that truly attempts to unify Transformer and Mamba in sequence level.
@@ A.3.2 Experiment Setup Details
  • PhoneBook is introduced in (Waleffe et al., 2024) and aims to evaluate the exact phone number of a specific person given a phone book of multiple people. There are two ways to construct a specific phone book: conventional construction and reverse construction. We use the open source tool (Faraglia & Other Contributors, ) to construct a completely random test set PhoneBook.

### v2

# TransMamba: A Sequence-Level Hybrid Transformer-Mamba Language Model

@@ Abstract
Transformers are the cornerstone of modern large language models, but their quadratic computational complexity limits efficiency in long-sequence processing. Recent advancements in Mamba, a state space model (SSM) with linear complexity, offer promising efficiency gains but suffer from unstable contextual learning and multitask generalization. Some works conduct layer-level hybrid structures that combine Transformer and Mamba layers, aiming to make full use of both advantages. This paper proposes TransMamba, a novel sequence-level hybrid framework that unifies Transformer and Mamba through shared parameter matrices (QKV and CBx), and thus could dynamically switch between attention and SSM mechanisms at different token lengths and layers. We design the Memory Converter to bridge Transformer and Mamba by converting attention outputs into SSM-compatible states, ensuring seamless information flow at TransPoints where the transformation happens. The TransPoint scheduling is also thoroughly explored for balancing effectiveness and efficiency. We conducted extensive experiments demonstrating that TransMamba achieves superior training efficiency and performance compared to single and hybrid baselines, and validated the deeper consistency between Transformer and Mamba paradigms at sequence level, offering a scalable solution for next-generation language modeling. Code and data are available at https://github.com/Yixing-Li/TransMamba.
@@ 1 Introduction
Transformers (Vaswani et al., 2017; Achiam et al., 2023; Touvron et al., 2023) are the foundation and mainstream model of modern deep learning (Zhao et al., 2023), showing dominating power in language modeling. Recently, Mamba has emerged (Gu and Dao, 2023) and been verified in various fields. Compared with Transformer, Mamba has linear computational complexity, high efficiency in processing long sequences, and lower training and inference costs (Qu et al., 2024). Nevertheless, its contextual learning and multi-task generalization capabilities are unstable (Waleffe et al., 2024). Transformer and Mamba have their own strengths and complement each other.
However, Transformer and Mamba have their own flaws that cannot be addressed by naive layer-shared hybrid structures (Yuan et al., 2024; Yang et al., 2024). For example, Transformer has faster training for short contexts while Mamba has better efficiency in longer contexts (see Table 2). Moreover, the naive static hybrid model has structural restrictions such as the order of Mamba and Transformer, mandatory ratios, etc (Lieber et al., 2024; Dao and Gu, 2024). The performance of the Hybrid model will deteriorate if these specific rules are not met, which greatly limits the exploration and breakthrough of the model.
Recently, Mamba2 (Dao and Gu, 2024) further enhances the performance of Mamba series, which reveals the surprising consistency of the attention of Transformer and the State Space Model (SSM) of Mamba. Furthermore, (Wang et al., 2024) performed distillation between Mamba and Transformer, distilling the QKV parameters of Attention to obtain CBx of SSM, verifying that the parameters can be interactively transferred as shown in Table 2. These motivate us that we can bravely utilize a set of shared parameters of QKV and CBx to build a joint Transformer-Mamba framework, which could flexibly decide which structure is suitable for the current training/inference in different layers/token lengths, taking advantage of both structures to balance effectiveness and efficiency while ensuring structural flexibility. Intuitively, to obtain the efficiency advantages of both structures, we can make the model adopt the Transformer mechanism for training on relatively short contexts and the SSM mechanism on long contexts. As shown in Figure 3, such prototype framework has only one set of parameters to flexibly switch between Transformer and Mamba for LM. In the first N tokens of the sequence, the parameter matrix is calculated using the attention mechanism. At a specific node in the sequence (which we call TransPoints from Transformer mode to Mamba2 mode), the parameter matrix is converted to the SSM mechanism for subsequent sequence generation, so as to achieve better training efficiency with better performance in sequences of different lengths.
@@ 2.1.1 Basic Notions and Consistency of Attention and SSM
We use the classic notation from the Transformer and Mamba papers. $\mathbf{QKV}$ denotes the key parameters (query, key, value) of Attention, and $\mathbf{L}$ denotes the additional mask matrix. $\mathbf{CBx}$ represents the key parameters in the SSM, $\Delta$ is used to control the discrete step size in SSM, and $\mathbf{A}$ is used to describe the global dependencies of the hidden state, which is similar to the mask matrix in Attention (Gu and Dao, 2023). In order to satisfy the classic symbolic representation and clear expression, we use $\mathbf{H}$ to denote the input embeddings for attention and SSM. The corresponding calculations are shown in Table 2.
Mamba2 (Dao and Gu, 2024) compares the underlying mechanisms of Transformer and Mamba, and introduces the dual form of SSM to illustrate the consistency between the two. In Table 2, we can find that the core mechanisms of Transformer and Mamba (attention and SSM) are completely symmetrical. Wang et al. (2024) aligned the QKV of the transformer weights with CBx of Mamba and performed distillation, achieving improved results on chat and long-text benchmarks. This once again shows that the core weights of Transformer and Mamba are transferable and unified. The above theories and research inspired us to build a bolder framework of TransMamba with a unified architecture of Transformer and Mamba.
@@ 2.1.2 Efficiency of Attention and SSM with different token lengths
As shown in Table 2, recent work (Dao and Gu, 2024) theoretically summarizes the FLOPs of Attention and SSM. $\mathrm{T}$ denotes the sequence length, $\mathrm{N}$ denotes the state dimension and $\mathrm{P}$ denotes the TransPoint value. When T is greater than N, Transformer has an advantage in efficiency on shorter sequences, while Mamba is efficient at training on long sequences due to its linear complexity of T. This advantage of Mamba’s training efficiency on longer contexts is present with most of the commonly-used model sizes, which also forms the motivation of our TransMamba that attempts to unleash the maximum potential of the flexible hybrid Transformer-Mamba structure in terms of effectiveness and efficiency.
@@ 2.4.2 Detailed TransPoint Schedule Designing
TransPoint schedule from token length aspect. Due to the flexibility of TransPoint scheduling at different layers and token lengths, the model structure of TransMamba varies at different positions. Suppose the number of layers of the model is L{L} and the length of the sequence is T{T}, there are LT{L}*{T} possible TransPoint schedules for our TransMamba with fixed parameters. We denote the value of TransPoint as P{P} (indicating that the tokens before position P{P} are modeled via Transformer and those after P{P} are encoded via Mamba), and we have the FLOPs for a TransMamba layer as follows:
@@
TransPoint schedule from layer aspect. We find that simply setting a global Transpoint for all layers (e.g., simultaneously at length 4,096) will result in unsatisfactory performance due to the sudden switching. To achieve better results, we need to set more diverse TransPoints at the layer level, which could gradually guide the model structure from pure Transformer to Mamba differently at various layers. To enable a smoother transformation, the TransPoints are placed separately but as close as possible to the optimal efficient $P$ according to Eq. 10. Specifically, our TransPoints cycle is performed every 8 layers, referring to the work (Dao and Gu, 2024) on hybrid structure. The mean of TransPoints is set slightly smaller than the optimal efficient point to enable more Mamba layers for better performance. TransPoints gradually transition from the beginning to the end of the sequence in a logarithmic trend (i.e., 0, 128, 256, 512, 1024, 2048, 4096, 8192), ensuring dispersion and smoothness.
@@ 3.1 Experiment Setup
We developed three baseline model families with various sizes (400M, 1.5B): Transformer (Shoeybi et al., 2019), Mamba2 (Dao and Gu, 2024), and Hybrid (Lieber et al., 2024). All models are developed based on the Megatron-LM library. Specifically, all models are unified in model size for fair comparisons. The models are pre-trained utilizing collected in-house dataset which consists of a cleaned combination of Chinese and English datasets. We trained all models on 83 billion tokens for all models. For evaluation, we aim to achieve robust conclusions across diverse domains. We conducted comprehensive evaluations involving 8 English tasks, including ARC-E, ARC-C (Clark et al., 2018), CoQA (Reddy et al., 2019), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), PhoneBook (Waleffe et al., 2024), BoolQ (Clark et al., 2019), LongBench-v2 (Bai et al., 2024). More details are shown in Appendix A.3.1 and A.3.2.
@@ 4 Related Works
Transformer has always been the focus of language model research (Beltagy et al., 2020; Liu et al., 2021; Tang et al., 2024), but its limitations in processing long sequences (Zhou et al., 2021; Behrouz et al., 2024) and the memory pressure caused by KV cache (Wang et al., 2020; Dao et al., 2022) are also difficult to solve. Mamba has the advantage of linear complexity based on the state space model (Gu and Dao, 2023; Zhang et al., 2024), but struggles in modeling complex contexts (Xiao et al., 2024).
Hybrid models (Chen et al., 2024; Lou et al., 2024; Ren et al., 2025) that combine the two are emerging, but most of the work simply cascades them (Hatamizadeh and Kautz, 2024; Lieber et al., 2024). Recent work (Dao and Gu, 2024; Han et al., 2024; Wang et al., 2024) has revealed the consistency of the underlying mathematics between them. However, there is no work that truly attempts to unify Transformer and Mamba in sequence level.
@@ A.3.2 Experiment Setup Details
  • PhoneBook is introduced in (Waleffe et al., 2024) and aims to evaluate the exact phone number of a specific person given a phone book of multiple people. There are two ways to construct a specific phone book: conventional construction and reverse construction. We use the open source tool (Faraglia and Other Contributors, ) to construct a completely random test set PhoneBook.